# Mirror symmetry and rigid structures of generalized K3 surfaces

### v5

@@ 4.7. Comments on relevant works
A generalized Calabi–Yau variety is a Fano variety $X$ whose derived category $\mathrm{D}^{b}(X)$ admits a semi-orthogonal decomposition and a component of it is Calabi–Yau category $\mathcal{A}_{X}$, which provides a deformation of the derived category of a Calabi–Yau variety of lower dimension. In this sense, $\mathcal{A}_{X}$ is thought to be a non-commutative Calabi–Yau variety.
Therefore, in dimension 2, there are two notions of non-commutative K3 surfaces: $\mathcal{A}_{X}$ for a Fano 4-fold $X$ and the derived category $\mathrm{D}^{b}(S,\alpha)$ of twisted coherent sheaves on a K3 surface $S$ (cf. Section 4.5). The relation between $\mathcal{A}_{X}$ and $\mathrm{D}^{b}(S,\alpha)$ has recently been elucidated and a certain correspondence is given in [1, 7]. However, there are $\mathcal{A}_{X}$ that do not correspond to $\mathrm{D}^{b}(S,\alpha)$ (and there are generalized Calabi–Yau structures that do not correspond to $\mathcal{A}_{X}$).
In [9] Sheridan and Smith proved among others homological mirror symmetry for some pairs of a rigid Calabi–Yau manifold and a generalized Calabi–Yau variety. They in particular discussed homological mirror symmetry for a rigid K3 surface and a cubic fourfold $X$ (op.cit. Section 1.7.2). The corresponding $\mathcal{A}_{X}$ is “non-geometric” and not in the above correspondence. Nevertheless it should correspond to a generalized Calabi–Yau structure and be compatible with our result. Note that our formulation is intrinsic and works for all singular K3 surfaces while [9] is concerned with (generalized) Calabi–Yau varieties obtained by the Batyrev-Borisov construction (but in all dimensions).

### v6

@@ 4.7. Comments on relevant works
A generalized Calabi–Yau variety is a Fano variety $X$ whose derived category $\mathrm{D}^{b}(X)$ admits a semi-orthogonal decomposition, with one component being the Calabi–Yau category $\mathcal{A}_{X}$. This category provides a deformation of the derived category of a Calabi–Yau manifold of lower dimension. In this sense, $\mathcal{A}_{X}$ is thought to be a non-commutative Calabi–Yau manifold.
Another notion of a non-commutative Calabi–Yau manifold arises from a certain algebra introduced in [15]; Here, the quotient category $\mathrm{qgr}(A/(f)):=\mathrm{gr}(A/(f))/\mathrm{tor}(A/(f))$ is considered, where $A$ is a specific choice of skew polynomial algebra and $f$ is a Fermat type element in the center of $A$. This construction yields a non-commutative projective scheme in the sense of Artin and Zhang [2].
In dimension 2, there are thus at least three notions of non-commutative K3 surfaces: $\mathcal{A}_{X}$ for a Fano 4-fold $X$, the quotient category $\mathrm{qgr}(A/(f))$ of modules of a certain algebra, and the derived category $\mathrm{D}^{b}(S,\alpha)$ of twisted coherent sheaves on a K3 surface $S$ (Remark 4.9). The relation between $\mathcal{A}_{X}$ and $\mathrm{D}^{b}(S,\alpha)$ has recently been elucidated, with a certain correspondence provided in [1, 14]. However, there exist instances of $\mathcal{A}_{X}$ that do not correspond to $\mathrm{D}^{b}(S,\alpha)$, and similarly, generalized Calabi–Yau structures that do not correspond to $\mathcal{A}_{X}$. The connection between the quotient category $\mathrm{qgr}(A/(f))$ and the other two cases remains unclear at this point.
Remark 4.11.
The author has been informed by Belmans that the methods in [5] should generalize to higher dimensions, and show that the derived category $\mathrm{D}^{b}(\mathrm{qgr}(A/(f))$ of modules of the non-commutative K3 surfaces arising in [15] is of the form $\mathrm{D}^{b}(S,\alpha)$. This work is currently in progress by the authors of [5].
In [18] Sheridan and Smith proved, among other results, homological mirror symmetry for certain pairs of a rigid Calabi–Yau manifold and a generalized Calabi–Yau variety. They in particular discussed homological mirror symmetry for a rigid K3 surface and a cubic fourfold $X$. The corresponding $\mathcal{A}_{X}$ is considered “non-geometric” and does not fall within the aforementioned correspondence. Nevertheless it should correspond to a generalized Calabi–Yau structure and align with our result. It is worth noting that our formulation is intrinsic and applicable to all singular K3 surfaces, whereas [18] focuses on (generalized) Calabi–Yau varieties obtained through the Batyrev-Borisov construction (in all dimensions).